1.true or false:
-3x-6y=-21
9x+6y=-9


2.which is addition method and substitution method?
x=7y+14
2x+15y=-75
4x-7y=14
2x+15y=-75
thank you so much

Answer by tutorcecilia(2144)

1.true or false:
-3x-6y=-21
9x+6y=-9
___________ [Subtract the equations]
6x=-30
x=-5 [Substitute and solve for y]
.

-3(-5)-6y=-21
15-6y=-21
-6y=-21-15
-6y=-36
y=6
.
checking:
9(-5)+6(6)=-9
-45+36 =-9
-9=-9, so the answer is true, x=-5 and y=6 for both equations.